What is the mass of a gold bar that is 7.379 × 10–4 m3 in volume? the density of gold is 19.3 g/cm3?
1
Expert's answer
2017-09-08T15:18:07-0400
p=m/v
m=p*v
m=737,9*19,3=14241g=14,241kg